Jackie invests $25,000, part as a 5.5% annual interest and the balance at 8.3% annual interest. How much is invested at each rate if Jackie receives a 1-year interest payment of $1571?
Answer by jorel1380(3719)

.055m+.083(25000-m)=1571
55m+83(25000-m)=1571000
55m+2075000-83m=1571000
2075000-1571000=83m-55m=28m
28m=504000
m=18000
Jackie invested $18000 at 5.5% interest and $7000 at 8.3% interest..